OpenStudy (sedatefrog712):

The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 82 feet with an initial velocity of 60 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. (2 points) Part B: What is the maximum height that the projectile will reach? Show your work. (2 points)

OpenStudy (sedatefrog712):

Part C: Another object moves in the air along the path of g(t) = 10 + 63.8t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know? (2 points)

OpenStudy (sedatefrog712):

@triciaal

OpenStudy (triciaal):

part A using the function H(t) s = 82 , v = 60 H(t) = 0 find t you will have a quadratic 2 values for t one value will be 0

OpenStudy (triciaal):

sorry part A just the equation Part B solve the equation

OpenStudy (triciaal):

Part B the time/2 for the maximum height reached

OpenStudy (sedatefrog712):

how would i set up the equation for part a

OpenStudy (triciaal):

part C Use a table to find the approximate solution to the equation H(t) = g(t) for any value of t find the value of Ht) and the value g(t) when H(t) = g(t) this is the time they will meet

OpenStudy (triciaal):

part a you are just substituting the given values v = 60 s = 82 etc

OpenStudy (triciaal):

part A is to write the equation

OpenStudy (sedatefrog712):

so it would look like this H(t) = -16t2 + 60t + 82

OpenStudy (triciaal):

yes

OpenStudy (sedatefrog712):

ok can you help me with part b also cause im bad at solving these

OpenStudy (triciaal):

for part A equation = 0 the ground has a height of 0 H(t) = 0

OpenStudy (sedatefrog712):

so its 0 = -16t2 + 60t + 82

OpenStudy (triciaal):

yes then solve to find t start (t = 0) and after how long will be back on ground (t= value when solved)

OpenStudy (sedatefrog712):

ok wait i thought part b where we needed to solve it

OpenStudy (sedatefrog712):

@triciaal

OpenStudy (triciaal):

yes Part A write the equation Part B solve the equation this will give you t find t/2 use this value to find the maximum height

OpenStudy (sedatefrog712):

ok can you help me with part b

OpenStudy (triciaal):

you need to give me time to read, interpret and do as well as answer any questions you have in-between

OpenStudy (sedatefrog712):

mkay im sorry for rushing

OpenStudy (triciaal):

now i forgot where i was

OpenStudy (triciaal):

let's solve the quadratic. you can divide everything by 4 to work with smaller numbers what is -16 t^2 + 60 t + 82 = 0 divided by 4?

OpenStudy (sedatefrog712):

-4t^2 + 15t + 20.5

OpenStudy (triciaal):

you are correct I thought it was 84

OpenStudy (triciaal):

let's use the formula to solve a = -4 b = 15 c = 20.5

OpenStudy (triciaal):